Both roots of the quadratic equation x^2-63x+k=0 are prime numbers.The number of possible value of k is

Grade:10

1 Answers

Vikas TU
7 years ago
Sum of Roots =>
a + b =  63
Product of roots = k = ab
now sum of roots is odd hece one root would be obvious even and that is even prime number 2.
hence a = 2 and b = 63 – 2 = 61
k = 61 * 2 = > 122